Required information [The following information applies to the questions displayed below.] On January 1 of this year, Olive C1. complete the amortization schedule.

2. When the bonds mature at the end of Year 4, what amount of principal will Olive pay investors?

3.  How much cash will be disbursed for interest each period and in total over the life of the bonds?

4. What was the annual market rate of interest on the date the bonds were issued?

5. What amount of interest expense will be reported on the income statement for Year 2 and Year 3?

6. What amount for bonds payable will be reported on the balance sheet at the end of Year 2 and Year 3?
Answer #1

1. Annual Interest Rate = 5200/52000 = 10%, Market Rate = 5019/52833 = 9.5%

Date Cash (10%) Interest Expense (9.5%) Amortization Balance
January 1, Year 1 52,833
End of Year 1 5,200                 5,019 181 52,652
End of Year 2

5,200

(52000*10%)

                5,002

(52652 *9.5%)

198

(5200-5002)

52,454
End of Year 3

5,200

(52000*10%)

                4,983

(52454*9.5%)

217

52,237

(52454-217)

End of Year 4

5,200

(52000*10%)

                4,963

237

(5200-4963)

52,000

2.

$52,000 will be the amount of principal which will be paid to investors.

3.

$5,200 will be disbursed each period and total interest disbursed will be $20,800

4.

Annual market rate of interest on the date the bonds were issued was 9.5% calculated as 5019/52833

5.

Interest expense reported in income statement in year 2 is $5,002 and Year 3 is $4983

6.

Amount reported in balance sheet in year 2 is $52,454and Year 3 is $52,237